All network equipment in Wheelock Student Center was replaced. Cabletron equipment was replaced with Cisco equipment.
Monthly Archives: December 2003
Google Search Appliance Upgraded
The Google search appliance was upgraded to operating system version 1.17.
The Search software was upgraded to version 3.4.14
Cbord Odyssey unresponsive to registers
The Cbord server reported to not be responding to inquiries from the cash registers at approximately 12:30 PST. Upon inspection of the server, it was discovered that pcAnywhere was hung waiting for a connection. We received a VFEP error when tring to access the Odyssey Control panel.
It is our belief that C-Bord support encountered errors when doing routine maintenance that hung the server.
RESOLUTION: We rebooted the Odyssey server and had the cash registers re-inquire.
Corporate Time slow
Serveral users have recently reported slow response from the Corporate Time server. During the course of our investigation of the problem, several orphaned ssh process were found on Big Ben and the log files were found to be quite large.
The orphaned processes were dealt with at 16:00 PST. The Corporate Time service was stopped at approximately 20:30 PST. The log files were rotated and the service restarted. The University web server was restarted to allow users to access the Corporate Time web interface.
Proxy Server Hang
The University’s proxy server hung this morning. the server was restarted, and returned to normal operation.
Internet Connection Instability II
The University’s internet connection was inoperable this morning from about 7 AM until about 10:30 AM. The problem was isolated to the PIX firewall. A number of logged outbound connection attemps (ca. 97000) from forged class A source addresses to a destination in Northern California were discovered. The sheer number of such attempts constituted a denial of service situation for the firewall.
The true source of the connection attempts was traced to a student computer in Todd/Phibbs. The students network port was inactivated, and the firewall was rebooted. This restored internet service.
Internet connection instability I
The University’s Internet connection was unstable from 7 PM to 9:30 PM. We are unsure of the reason. Checked for unusual traffic at the core router, and found none. The traffic analysis graph shows an unusually flat usage level at about 5 Mbps during this the period 7PM to 9PM, then usage dropped to almost 0, and is now recovering.
The internet service provider has been called.
The disk partitions on Camano were modified to isolate disk space for BJP, provide more swap space, and more space for Oracle. Work completed before the 9am end of PM.
Cascade web (Enterprise Manager) was restarted via command line, because the oracle account .profile file was incorrect.
BJP was started by command line, but issues with the startup procedure required a few extra steps to be taken. BJP was up and running as of 9am.
The HP laserjet 9000 (syslaser1) was tweaked to try and determine source of multiple print jobs. Items changed on printer:
Clearable Warnings (changed from Job to On)
Auto Continue (changed from On to Off)
PIX Firewall reconfigured
The number of ports used by the ezproxy server was extended to 5000-7000, to allow more sessions.